### v1

# QCD phase transition at finite temperature and chemical potential with the non-extensive statistics

Abstract
With the application of the non-extensive correction to the equation of state (EoS) in the parton (hadron resonance) gas at high (low) temperature and a smooth interpolation of these two equation of states, a non-extensive version of EoS is constructed to investigate the properties of quantum chromodynamics (QCD) phase transition at finite temperature ($T$) and chemical potential ($\mu$). It is based on non-extensive statistics and characterized by a non-extensivity parameter $q$, which measures the degrees of deviation from the thermal equilibrium. We explore the QCD phase transition with $\mu=0$ MeV in the chiral and non-chiral limit cases, as well as the transition with finite $\mu$ in the non-chiral limit case. It is found that the dimensionless thermodynamic quantities such as the entropy density, the pressure, the energy density, the specific heat at constant volume and the trace anomaly are sensitive to the deviation of $q$ from unity and they become large both in the hadronic and quark-gluon plasma (QGP) phases with the increase of $q$. In the chiral limit case, the sound velocity squared ($(c_{s}^{2})_{q}$) is insensitive to the deviation. However, in the non-chiral limit case, this deviation leads to nontrivial corrections of $(c_{s}^{2})_{q}$ in the vicinity of the critical point ($T_{c}$) and at lower temperatures. Moreover, these dimensionless thermodynamic quantities are sensitive to the deviation of $\mu$ from zero and they become large both in the hadronic and QGP phases with the increase of $\mu$. However, for $(c_{s}^{2})_{q}$, it increases in the vicinity of $T_{c}$ and decreases at lower temperature. Finally, we observe that our results in the non-chiral limit case with $q=1$ agree well with those from the Lattice QCD, the hadron resonance gas model, and the Thermal-Fist fit to the hadron yields in high energy nuclear collisions in the low temperature region up to $T\sim 150$ MeV.
I Introduction
Quantum-Chromodynamics (QCD) predicts that at high temperatures and energy densities there exists a deconfining transition of the matter governed by strong interactions to a new state denoted as the quark-gluon plasma (QGP). QGP is expected to be produced in ultra-relativistic heavy-ion collisions. Exploring the properties of QCD phase transition is one of the main goals in heavy-ion collisions.
In recent years, several effective models, such as the bag model [1], the Nambu-Jona-Lasinio (NJL) model [2, 3], the linear sigma model [4], the quark-meson model [5, 6], the Dyson-Schwinger equations approach [7, 8], the functional renormalization group approach [9, 10, 11], the hadron resonance gas (HRG) and statistical models [12, 13, 14, 15], and the Lattice QCD (LQCD) [14, 15, 16, 17], have been utilized to investigate the properties of the QCD phase transition. In these models, a thermal equilibrium is assumed and the extensive statistical approach, such as the Boltzmann-Gibbs (BG) statistics, the Bose-Einstein statistics or the Fermi-Dirac statistics is applied.
However, such an approach is only valid when the heat bath of the system is homogeneous. In reality, this condition cannot always be satisfied. For example, in high-energy nucleus-nucleus (AA) collisions, extreme conditions of density and temperature give rise to strong intrinsic fluctuations and long-range correlations [18, 19, 20, 21]. Thus, instead of an equilibrium state, some kind of stationary state near the equilibrium ($q$-equilibrium) is expected to be formed [22, 23]. This phenomenon can be dealt with non-extensive statistics (Tsallis statistics) [24] where the degree of deviation from the equilibrium is described by a single parameter $q$.
In recent years, Tsallis statistics has been widely employed to describe the multi-particle production processes in high-energy proton-proton (pp) [25, 26, 27, 28, 29, 30, 31, 32, 33] and AA [34, 35, 36, 37, 38, 39, 40, 41, 42] collisions. In most of these applications, the Tsallis distribution or the Tsallis-extended blast-wave model is simply fitted to the particle transverse momentum ($p_{\rm T}$) spectra. The extracted values of $q$ from the fits are usually larger than 1. The deviation of $q$ from unity represents the intrinsic fluctuations of the temperature [18] or of the mean value of the charged-particle multiplicity [43] in the hadronizing system. In case of high-energy AA collisions, the fluid dynamics is not considered in these applications, which could affect the accuracy of the results. In light of such a situation, (0+1)-, (1+1)-, (2+1)- and (3+1)-dimensional hydrodynamic models based on non-extensive statistics were developed as a description of relativistic heavy-ion collisions [44, 45, 46, 47]. This is a reasonable extension of the hydrodynamic model, as far-from-equilibrium evolution can still obey the hydrodynamic equations of motion, provided that the system is locally isotropized in the co-moving frame of the fluid [48, 49]. It was shown that the non-extensive hydrodynamic model can describe the charged hadron $p_{\rm T}$ spectra up to 6-8 GeV/c, which is larger than the typical validity range of the extensive hydrodynamic description, 2-3 GeV/c. Moreover, the elliptic flow of charged particles was reduced with the introduction of Tsallis statistics to the hydrodynamic model. The Tsallis statistics has also been applied on the nuclear equation of state (EoS) and the QGP EoS in the framework of a relativistic mean-field approach and the bag model, respectively. In the former model, the interactions of nucleons are mediated by the exchange of isoscalar and isovector mesons, $\sigma$, $\omega$ and $\rho$ [50, 51, 52, 53]. In the latter model, the quark matter is described as a gas of free quarks and the non-perturbative effects are simulated by the bag constant $B$ [50, 51, 53]. It was observed that both the nuclear EoS and the QGP EoS became stiffer when the non-extensive statistic effect was considered. Furthermore, the Tsallis statistics has been taken into account in the NJL model to describe the QCD phase transition [54, 55, 56, 57]. It was found that with the increase of $q$ the critical end point moved toward the direction of larger chemical potential and lower temperature.
In this study, we try to construct a non-extensive version of EoS to investigate the generic feature of the QCD phase transition at finite temperature and chemical potential. We first explore the QCD phase transition with zero chemical potential in the chiral and non-chiral limit cases. In the former case, the dominant excitations in the hadronic phase are the massless pions, while in the QGP phase are massless $u$, $d$ quarks and gluons. In the latter case, in the QGP phase the masses of $u$, $d$ and $s$ quarks are considered, and in the hadronic phase the hadrons and hadron resonances composed of $u$, $d$ and $s$ quarks with masses below $2.5$ GeV$/c^{2}$ are considered. Then we investigate the QCD phase transition with finite chemical potential in the non-chiral limit case. We will investigate the effect of the chemical potential and the non-extensive statistics on the temperature dependence of the dimensionless thermodynamic quantities such as the entropy density, the pressure, the energy density, the trace anomaly, the heat capacity at the constant volume, and the sound velocity squared. Moreover, we will compare our results with those from the HRG, the LQCD, and the fit of the experimental yields of hadrons at the LHC and RHIC energies using the Thermal-Fist package [58].
This paper is organized as follows. In sect. II, we will describe the Tsallis statistics and construct a non-extensive version of EoS by applying the non-extensive correction to the EoS in the parton (hadron resonance) gas at high (low) temperature and making a smooth interpolation of these two equation of states. In sect. III, we will discuss the effect of the non-extensive statistics and the chemical potential on the thermodynamic quantities and compare our results with different models. Finally, the conclusion is given in sect. IV.
II The model
II.1 Tsallis statistics
The non-extensive statistics was first proposed by Tsallis [59]. It generalizes the BG statistics and defines the entropy of a system as
$S_{q}=k\left(1-\sum\limits_{\begin{subarray}{c}R=1\end{subarray}}^{W}p_{R}^{q}\right)/(q-1),$ (1)
where $k$ is a positive constant; $W\in N$ is the total number of microscopic configurations of the system; $p_{R}$ is the probability associated with the $R^{th}$ state and satisfies $\sum_{R=1}^{W}p_{R}=1$; $q$ characterizes the degree of non-extensivity reflected in the following pseudo-additive rule for a composite system consisting of two individual systems $A$ and $B$,
$\frac{S_{q}(A+B)}{k}=\frac{S_{q}(A)}{k}+\frac{S_{q}(B)}{k}+(1-q)\frac{S_{q}(A)}{k}\frac{S_{q}(B)}{k}.$ (2)
In the limit $q\rightarrow 1$, the expression in Eq. (1) approaches the BG entropy,
$\lim\limits_{\begin{subarray}{c}q\end{subarray}\rightarrow 1}S_{q}=-k\sum_{R=1}^{W}p_{R}\textrm{ln}p_{R}.$ (3)
In order to determine the probability distribution of a state $R$ in a grand canonical system, we need to maximize the entropy under the internal energy and particle number constraints,
$\bar{E}=\sum\limits_{\begin{subarray}{c}R=1\end{subarray}}^{W}p_{R}^{q}E_{R},$ (4)
$\bar{N}=\sum\limits_{\begin{subarray}{c}R=1\end{subarray}}^{W}p_{R}^{q}N_{R},$ (5)
where $\bar{E}$ and $\bar{N}$ are, respectively, the average energy and the average number of particles. The maximization of the entropy under constraints in Eqs. (4) and (5) leads to the variational equation,
$\small\frac{\delta}{\delta p_{R}}\left(\frac{S_{q}}{k}-\alpha\sum_{R=1}^{W}p_{R}-\beta\sum_{R=1}^{W}p_{R}^{q}E_{R}-\gamma\sum_{R=1}^{W}p_{R}^{q}N_{R}\right)=0,$ (6)
where $\alpha$, $\beta$ and $\gamma$ are the Lagrange multipliers associated with the total probability, the total energy and the total number of particles, respectively. Following the standard thermodynamic definitions, such as $\beta=1/T$ and $\gamma=-\beta\mu$ with $T$ and $\mu$, respectively, being the temperature and the chemical potential of the system, the probability distribution of the state $R$ in the grand-canonical ensemble is written as
$p_{R}=[1+(q-1)(E_{R}-\mu N_{R})/T]^{1/(1-q)}/Z_{q},$ (7)
where
$Z_{q}=\sum_{R}[1+(q-1)(E_{R}-\mu N_{R})/T]^{1/(1-q)}$ (8)
is the partition function of the grand-canonical system.
The quantum state of the system is uniquely determined when occupation numbers of the one-particle states are provided. The total energy and the total particle numbers of the system in the state $R$ then are automatically given by
$E_{R}$ $=n_{1}\varepsilon_{1}+n_{2}\varepsilon_{2}+\cdots+n_{k}\varepsilon_{k}+\cdots,$ (9)
$N_{R}$ $=n_{1}+n_{2}+\cdots+n_{k}+\cdots,$ (10)
where $\varepsilon_{k}$ and $n_{k}$ are, respectively, the energy of a particle and the number of particles in the state $k$. Substituting Eqs. (9) and (10) into Eq. (8), one gets the partition function as
$Z_{q}$ $=\sum_{n_{1},n_{2}\cdots,n_{k},\cdots}\exp_{q}\left(-\frac{(\varepsilon_{1}-\mu)n_{1}}{T}\right)$ (11)
$\otimes_{q}\exp_{q}\left(-\frac{(\varepsilon_{2}-\mu)n_{2}}{T}\right)\otimes_{q}\cdots$
$\otimes_{q}\exp_{q}\left(-\frac{(\varepsilon_{k}-\mu)n_{k}}{T}\right)\otimes_{q}\cdots$
$=(\prod_{k=1}^{\infty})_{q}\sum_{n_{k=0}}^{\infty}\exp_{q}\left(-\frac{(\varepsilon_{k}-\mu)n_{k}}{T}\right),$
where $\textrm{exp}_{q}(x)=[1+(1-q)x]^{(1/(1-q))}$, $\textrm{exp}_{q}(x)\otimes_{q}\textrm{exp}_{q}(y)=\textrm{exp}_{q}(x+y)$, $(\prod)_{q}$ means the successive $q$-multiplication. When taking into account the $d$ internal degrees of freedom for the non-interacting bosons or fermions, the grand-canonical partition function becomes
$Z_{q}\approx(\prod_{k=1}^{\infty})_{q}\left[1\pm{\rm exp}_{q}\left(-\frac{\varepsilon_{k}-\mu}{T}\right)\right]^{\pm d},$ (12)
where the plus and minus signs, respectively, refer to fermions and bosons. The derivation of Eq. (12) is shown in the appendix A.
II.2 Equation of states in the hadronic and QGP phases
In the framework of extensive statistics, the grand-canonical potential $\Omega$ is related to the grand-canonical partition function $Z$ as follows [60]
$\Omega=-T\ln Z.$ (13)
In the non-extensive statistics, the relation between the grand-canonical potential $\Omega_{q}$ and the grand-canonical partition function $Z_{q}$ is similar to that in the extensive statistics,
$\Omega_{q}=-T\ln_{q}Z_{q},$ (14)
where the $q$-logarithm ${\rm ln}_{q}x$ is defined as [61]
$\ln_{q}x\equiv\frac{x^{1-q}-1}{1-q}.$ (15)
The derivation of Eq. (14) is presented in the appendix B.
It has been shown that in the non-extensive statistics the thermodynamic consistency is also satisfied [26, 20]. Thus, the pressure is given by
$P_{q}$ $=-\frac{\Omega_{q}}{V}$ (16)
$=\frac{T\ln_{q}Z_{q}}{V}$
$=\frac{T}{V}\sum_{k=1}^{\infty}\ln_{q}\left[1\pm{\rm exp}_{q}\left(-\frac{\varepsilon_{k}-\mu}{T}\right)\right]^{\pm d},$
where the identity $\ln_{q}(x\otimes_{q}y)=\ln_{q}x+\ln_{q}y$ is used. In the large volume limit, the summation in Eq. (16) is replaced by the integral in the 6-dimensional phase space,
$\sum_{k}\rightarrow V\int\frac{d^{3}k}{(2\pi)^{3}}.$ (17)
Therefore, the pressure is rewritten as
$P_{q}=T\int\frac{d^{3}k}{(2\pi)^{3}}\ln_{q}\left[1\pm{\rm exp}_{q}\left(-\frac{\varepsilon_{k}-\mu}{T}\right)\right]^{\pm d},$ (18)
where $\varepsilon_{k}=\sqrt{m^{2}+k^{2}}$, $m$ and $k$ are, respectively, the rest mass and the momentum of the particle.
The entropy density for a multi-component system with $N$ number of species then is expressed as
$s_{q}=\sum_{i=1}^{N}\partial P_{q}^{i}/\partial T,$ (19)
where $P_{q}^{i}$ is the pressure for the single-component system with the particle type $i$. One key factor in the calculation of the entropy density is the internal degrees of freedom $d$. In the chiral limit case, the dominant excitations in the hadronic (QGP) phase are the massless pions ($u$, $d$ quarks and gluons). The degeneracy factor for massless pions is $d_{\pi}=N_{f}^{2}-1$, where $N_{f}=$ 2 is the number of quark flavors considered. The degeneracy factors for the massless quarks and gluons are, respectively, $d_{g}=2_{\textrm{spin}}\times(N_{c}^{2}-1)=$ 16 and $d_{q}=2_{\textrm{spin}}\times 2_{q\bar{q}}\times N_{c}\times N_{f}=$ 24, where $N_{c}=3$ is the number of colors. For the non-chiral limit case, we consider the masses of $u$, $d$ and $s$ quarks (hadrons composed of these quarks) in the QGP (hadronic) phase. The degeneracy factors for quarks (antiquarks) with a given species and gluons are, respectively, $d_{q(\bar{q})}=6$ and $d_{g}=16$, while the degeneracy factor for hadrons with a given species is $d_{H}=2S+1$, where $S$ is the hadron spin. Another key factor in the calculation of the entropy density is the chemical potential. The chemical potential for $i^{th}$ particle species is $\mu_{i}=B_{i}\mu_{B}+S_{i}\mu_{S}+Q_{i}\mu_{Q}$, where $B_{i}$, $S_{i}$ and $Q_{i}$ are, respectively, the particle’s baryon, strangeness and electric charge numbers, $\mu_{B}$, $\mu_{S}$ and $\mu_{Q}$ are the respective chemical potentials of the system.
II.3 Smooth crossover around the critical temperature
In order to match the equation of states in the hadronic and QGP phases around the crossover temperature $T_{c}$, we make a smooth interpolation of $s_{q}(T)$ between the hadronic gas at low $T$ and the QGP at high $T$. A simplest possible parametrization of the entropy density $s_{q}(T)$ satisfying the thermodynamic inequality $\partial s_{q}(T)/\partial T\geq 0$ and the third law of thermodynamics $s_{q}(T\rightarrow 0)\rightarrow\textrm{constant}$ is [12]
$s_{q}(T)=f(T)s_{q}^{\textrm{H}}(T)+(1-f(T))s_{q}^{\textrm{QGP}}(T),$ (20)
where $f(T)$ is a weighting function,
$f(T)=\frac{1}{2}[1-{\rm tanh}((T-T_{c})/\it\Gamma)],$ (21)
with $\it\Gamma$ setting the width of the phase transition region. For $T$ satisfying $\lvert T-T_{c}\rvert>\it\Gamma$, $s_{q}(T)$ quickly approaches $s_{q}^{\textrm{H}}$ below $T_{c}$ and $s_{q}^{\textrm{QGP}}$ above $T_{c}$.
The pressure $P_{q}(T)$, the energy density $\varepsilon_{q}(T)$, the specific heat at constant volume $(C_{V})_{q}(T)$, and the sound velocity squared $(c_{s}^{2})_{q}(T)$ then can be, respectively, obtained using the following expressions,
[…]
$(C_{V})_{q}(T)=\frac{\partial\varepsilon_{q}(T)}{\partial T}\big{|}_{V}=T\frac{\partial s_{q}(T)}{\partial T}+\sum_{i}\mu_{i}\frac{\partial s_{q}(T)}{\partial\mu_{i}},$ (24)
$(c_{s}^{2})_{q}(T)=\frac{\partial P_{q}(T)}{\partial\varepsilon_{q}(T)}=\frac{s_{q}(T)}{(C_{V})_{q}(T)},$ (25)
where the index $i$ in the summation runs over all the particle species.
III Results and discussions
III.1 Zero chemical potential
For the system with zero chemical potential, we will first consider the QCD phase transition in the chiral limit case, where the dominant excitations in the hadronic (QGP) phase are the massless pions (massless $u$, $d$ quarks and gluons). In order to see the mass effect, we also consider the QCD phase transition in the non-chiral limit case. In this case, the $u$, $d$ and $s$ quarks in the QGP phase are considered, and their masses are, respectively, $m_{u}=3.503$ MeV$/c^{2}$, $m_{d}=3.503$ MeV$/c^{2}$, and $m_{s}=96.4$ MeV$/c^{2}$ [62]. In the hadronic phase, hadrons and hadron resonances composed of $u$, $d$ and $s$ quarks with masses below $2.5$ GeV$/c^{2}$ [16] from the Particle Data Group [63] are considered.
Figure 1: The upper (lower) panels: $s_{q}/T^{3}$ $(P_{q}/T^{4})$ as a function of $T/T_{c}$. The left (right) panels correspond to the chiral (non-chiral) limit case. The solid and dashed lines correspond to $q=$ 1.05 and $q=$ 1, respectively.
Figure 2: The upper (lower) panels: $\varepsilon_{q}/T^{4}$ $((C_{V})_{q}/T^{3})$ as a function of $T/T_{c}$. The left (right) panels correspond to the chiral (non-chiral) limit case. The solid and dashed lines correspond to $q=$ 1.05 and $q=$1, respectively.
Figure 3: The upper (lower) panels: $\Delta_{q}$ ($(c_{s}^{2})_{q}$) as a function of $T/T_{c}$. The left (right) panels correspond to the chiral (non-chiral) limit case. The solid and dashed black lines correspond to $q=$1.05 and $q=$ 1, respectively. In the lower left panel, the solid and dashed blue lines correspond to $m_{\pi}$= 140 ${\rm MeV}/c^{2}$.
Figures 1, 2 and 3, respectively, present the dimensionless thermodynamic quantities such as the entropy density $s_{q}/T^{3}$, the pressure $P_{q}/T^{4}$, the energy density $\varepsilon_{q}/T^{4}$, the specific heat at constant volume $(C_{V})_{q}/T^{3}$, the trace anomaly $\Delta_{q}=(\varepsilon_{q}-3P_{q})/T^{4}$ and the sound velocity squared $(c_{s}^{2})_{q}$ as a function of the temperature for two non-extensive parameters $q=$ 1 (the dashed line) and $q=$ 1.05 (the solid line). The left (right) panels of these figures show the temperature dependence of these thermodynamic quantities in the chiral (non-chiral) limit case. In these figures, $T_{c}$ is taken to be 156.5 MeV [64], $\it\Gamma$ is set as 0.05 $T_{c}$, which is identical to the choice made in Ref. [12]. The qualitative behavior of these thermodynamic quantities in both cases can be summarized as follows.
(i) There is a rapid increase of the entropy density in a narrow region of $T$ around $T_{c}$ (with a width $\sim$ 8 MeV). In the high temperature region, $s_{q}/T^{3}$ approaches the Stefan-Boltzmann (SB) limit, and the SB value in the chiral limit case is smaller than that in non-chiral limit case. This is due to reason that in the latter case one additional quark flavor is considered. In the low temperature region, $s_{q}/T^{3}$ approaches a constant non-zero value for the former case, while it approaches 0 for the latter case.
(ii) $P_{q}/T^{4}$ increases rather slowly with $T>T_{c}$ and approaches the SB limit at high temperature. As shown in Eq. (22), $P_{q}(T)$ is expressed as an integral of $s_{q}(T)$, thus such a continuous and slow rise is naturally expected [12].
(iii) There is a peak just above $T_{c}$ for the temperature dependence of $\varepsilon_{q}/T^{4}$. The existence of such a peak is due to the rapid increase of $s_{q}(T)$ and the slow rise of $P_{q}(T)$ above $T_{c}$ [12].
(iv) $(C_{V})_{q}/T^{3}$ has an obvious peak near $T_{c}$. The specific heat at constant volume is defined as the derivative of the energy density with respect to the temperature. The dimensionless $\varepsilon_{q}/T^{4}$ has a peak above $T_{c}$, which leads to the peak of $(C_{V})_{q}/T^{3}$ near $T_{c}$.
(v) $\Delta_{q}$ also has a prominent peak near $T_{c}$. Such a peak is a consequence of the rapid increase of $s_{q}(T)$. It vanishes when the temperature approaches 0 and $\infty$.
(vi) $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$ and $\Delta_{q}$ are sensitive to the deviation of $q$ from unity and they become large both in the hadronic and QGP phases with the increase of $q$. This can be interpreted as the effect that at a fixed temperature the particle yield in the non-extensive distribution with $q>$ 1 is always larger than that in the extensive one.
(vii) There is a sudden drop of $(c_{s}^{2})_{q}$ in the narrow region $|T-T_{c}|<\it\Gamma$. This could be explained as follows. $P_{q}(T)$ $(\varepsilon_{q}(T))$ varies slowly (rapidly) with temperature across the critical point, which leads to the softening of the equation of state in the critical region. In order to see the effect of the finite pion mass on $(c_{s}^{2})_{q}$ below $T_{c}$, the hadroinc sound velocity with $m_{\pi}=$ 140 ${\rm MeV}/c^{2}$ (blue solid and dashed lines) is also presented in the lower left panel of Fig. 3. In the chiral limit case, away from $T_{c}$, $(c_{s}^{2})_{q}$ approaches the relativistic limit, $1/3$. In the non-chiral limit case, $(c_{s}^{2})_{q}$ approaches $1/3$ in the high temperature region and 0 in the low temperature region. Moreover, in the former case, $(c_{s}^{2})_{q}$ is insensitive to the deviation of $q$ from unity. However, in the latter case, this deviation leads to nontrivial corrections of the sound velocity in the vicinity of the critical point and at lower temperatures.
III.2 Finite chemical potential
For the system with finite chemical potential, we only consider the QCD phase transition in the non-chiral limit case. The baryon, strangeness and electric charge potentials are, respectively, set as $\mu_{B}=$ 398.2 MeV, $\mu_{S}=$ 89.5 MeV and $\mu_{Q}=$ 0 MeV, which correspond to the chemical potentials of the system created in the central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7 GeV [65]. Figure 4 shows the temperature dependence of $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $\Delta_{q}$, $(C_{V})_{q}/T^{3}$ and $(c_{s}^{2})_{q}$. The solid and dashed black (blue) lines correspond to $q=$ 1.05 and $q=$ 1 with zero (finite) chemical potential, respectively. In the figure, $T_{c}$ is taken to be 156.5 MeV [64], $\it\Gamma$ is set as 0.05 $T_{c}$. The qualitative behaviors of the thermodynamic quantities are summarized as follows.
(i) $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$ and $\Delta_{q}$ are sensitive to the deviation of the chemical potential from zero and they become large both in the hadronic and QGP phases with the increase of $\mu$. This is due to the reason that at a given temperature a larger population is expected for the system with finite chemical potential. A similar increase was observed for the normalized pressure and trace anomaly obtained from the Dyson-Schwinger equations when the values of quark chemical potential were increased [66].
(ii) With the increase of the chemical potential, the sound velocity squared $(c_{s}^{2})_{q}$ increases in the vicinity of $T_{c}$ and decreases at lower temperature, which is different from the chemical potential dependence of $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$ and $\Delta_{q}$.
(iii) At high temperature, these thermodynamic quantities with finite chemical potential approach the respective ones with zero chemical potential. At low temperature, the thermodynamic quantities with both zero and finite chemical potential approach 0 when $T\rightarrow 0$.
(iv) As shown in the previous section, at zero chemical potential the thermodynamic quantities are sensitive to the deviation of $q$ from unity. With the consideration of both the effects of finite chemical potential and the non-extensivity, an obvious change of the behavior for the thermodynamic quantities is observed.
Figure 4: $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $\Delta_{q}$, $(C_{V})_{q}/T^{3}$ and $(c_{s}^{2})_{q}$ as a function of $T/T_{c}$ in the non-chiral limit case. The solid and dashed black lines correspond to $q=$ 1.05 and $q=$ 1 with zero chemical potential, respectively. The solid and dashed blue lines correspond to $q=$ 1.05 and $q=$ 1 with $\mu_{B}=$ 398.2 MeV, $\mu_{S}=$ 89.5 MeV and $\mu_{Q}=$ 0 MeV, respectively.
III.3 Comparison with other models
Figure 5: Upper panel: $P_{q}/T^{4}$ as a function of $T$ with zero chemical potential for $q=$ 1.05 (the solid line) and 1 (the dashed line) in the non-chiral limit case. The results from the HRG model, the LQCD and the Thermal-Fist fit are, respectively, shown as the dotted line, the solid black circles and the empty markers. Lower panel: $P_{q}/T^{4}$ as a function of $T$ with finite chemical potential for $q=$ 1.05 (the solid line) and $q=$ 1 (the dashed line) in the non-chiral limit case. The results from the HRG model and the Thermal-Fist fit are, respectively, shown as the dotted line and the empty markers.
The upper panel of Fig. 5 displays the comparison of the temperature dependence of $P_{q}/T^{4}$ calculated with zero chemical potential for different models. Our results with $q=$ 1.05 and $q=$ 1 in the non-chiral limit case are, respectively, presented as the solid and dashed lines. The result from the HRG model is shown as the dotted line, while the result from the LQCD is presented as the solid black circles. The dependence extracted from the fit of experimental yields of hadrons at the central centrality in pp collisions at $\sqrt{s}=$ 7 TeV [67, 68, 69] and 13 TeV [70, 71, 72], p-Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV [73, 74, 75], Pb-Pb collisions at $\sqrt{s_{\rm NN}}=$ 2.76 TeV [76, 77, 78, 79] and 5.02 TeV [80, 81], and Xe-Xe collisions at $\sqrt{s_{\rm NN}}=$ 5.44 TeV [82] using the Thermal-Fist package [58] are shown as empty markers. Thermal-Fist is a C++ package designed for the analyses of particle productions in relativistic heavy-ion collisions within the framework of the HRG model. In the package, there are several options regarding the ensemble with which to treat the baryon number, strangeness and electric charge. In this work, we choose the grand canonical ensemble, which is widely applied to heavy-ion collisions [83, 84, 85, 86]. It is shown that our result with $q=1$ agrees well with the HRG model and the LQCD data up to $T\sim$ 150 MeV. In our model, $P_{q}(T)$ quickly approaches the pressure at the hadronic phase with $T<T_{c}-\it\Gamma$. When $q\rightarrow 1$, the Tsallis statistics approaches the BG statistics valid for systems under the thermal equilibrium. Thus, the agreement at low temperature is reasonably expected as the HRG model and the LQCD calculations are based on equilibrium considerations. Moreover, our result with $q=1$ is consistent with the results from the Thermal-Fist fit to the hadron yields at the central centrality in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=$ 2.76 TeV (triangle-left) and 5.02 TeV (triangle-up), and in Xe-Xe collisions at $\sqrt{s_{\rm NN}}=$ 5.44 TeV (triangle-down). However, it overestimates the results from the Thermal-Fist fit to the hadron yields in p-Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV (triangle-right), and in pp collisions at $\sqrt{s}=$ 7 TeV (empty circle) and 13 TeV (empty square). In pp and p-Pb collisions, the temperature of the system is around 160 MeV, which is in the phase transition region $\lvert T-T_{c}\rvert<\it\Gamma$, leading to the overestimate of our model for the pressure. At high temperature, our result approaches the SB limit of the pressure for idea parton gas. However, in LQCD, partons interact with each other, which leads to the deviation from the limit. In the HRG model, only the hadron gas is considered and the phase transition is absent, thus the pressure approaches infinity at high temperature.
Table 1: Chemical potentials in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7, 11.5 and 200 GeV.
$\sqrt{s_{\rm NN}}$ (GeV) centrality $\mu_{B}$ (MeV) $\mu_{S}$ (MeV) $\mu_{Q}$ (MeV)
[…]
The lower panel of Fig. 5 shows the comparison of the temperature dependence of $P_{q}/T^{4}$ calculated with three sets of finite chemical potentials (see table 1) for different models. They correspond to the chemical potentials in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7, 11.5 and 200 GeV [65]. Our results with $q=$ 1.05 and $q=$ 1 in the non-chiral limit case are, respectively, presented as the solid and dashed lines. The results from the HRG model are shown as the dotted line. The dependence extracted from the Thermal-Fist fit of experimental yields of hadrons in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7, 11.5 and 200 GeV [65, 87, 88, 89] are shown as empty blue, red and black markers, respectively. It is found that our results with $q=1$ agree well with the HRG model up to $T\sim$ 150 MeV. Moreover, our results with $q=1$ are consistent with the results from the Thermal-Fist fit to the hadron yields in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 7.7 GeV (triangle-right) and 11.5 GeV (triangle-left). However, they overestimate the result from the Thermal-Fist fit to the hadron yields in central Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV (triangle-up). The temperature of the system in Au-Au collisions at $\sqrt{s_{\rm NN}}=$ 200 GeV is around 160 MeV, falling in the phase transition region and leading to this overestimate.
IV Conclusions
In this work, by applying the non-extensive correction to the EoS in the parton (hadron resonance) gas at high (low) temperature and making a smooth interpolation of these two equation of states, we have constructed a non-extensive version of EoS to investigate the properties of QCD phase transition with zero chemical potential in the chiral and non-chiral limit cases. We have also investigated the QCD phase transition with finite chemical potential in the non-chiral limit case. We find that the dimensionless quantities $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$ and $\Delta_{q}$ are sensitive to the deviation of $q$ from unity and they become large both in the hadronic and QGP phases with the increase of $q$. In the chiral limit case, $(c_{s}^{2})_{q}$ is insensitive to the deviation of $q$ from unity. However, in the non-chiral limit case, this deviation leads to nontrivial corrections of $(c_{s}^{2})_{q}$ in the vicinity of the critical point and at lower temperatures. Moreover, these dimensionless thermodynamic quantities are sensitive to the deviation of $\mu$ from zero and they become large both in the hadronic and QGP phases with the increase of $\mu$. However, $(c_{s}^{2})_{q}$ increases in the vicinity of the phase transition temperature and decreases at lower temperature. Finally, in the non-chiral limit case, we observe that our results with $q=1$ agree well with the LQCD, the HRG and the results form the Thermal-Fist fit to the hadron yield in high energy nuclear collisions in the low temperature region up to $T\sim 150$ MeV. These findings will deepen our understanding for the properties of the QCD phase transition at finite temperature and chemical potential both in the extensive and non-extensive statistics.
Acknowledgements.
This work is supported by the Scientific Research Foundation for the Returned Overseas Chinese Scholars, State Education Ministry, by Natural Science Basic Research Plan in Shaanxi Province of China (program No. 2023-JC-YB-012), and by the National Natural Science Foundation of China under Grant Nos. 11447024, 11505108 and 12275204.
Appendix A: The derivation of Eq. (12)
In the framework of the non-extensive statistics, the grand-canonical partition function of particles with $d$ internal degrees of freedom is given by
$Z_{q}=(\prod_{k=1}^{\infty})_{q}\left[\sum_{n_{k}=0}^{\infty}\exp_{q}\left(-\frac{(\varepsilon_{k}-\mu)n_{k}}{T}\right)\right]^{d},$ (A1)
where $\varepsilon_{k}$ and $n_{k}$ are, respectively, the energy of a particle and the number of particles in the state $k$; $T$ and $\mu$ are, respectively, the temperature and the chemical potential of the system.
For non-interacting fermions, their grand-canonical partition function is written as
$Z_{q}^{F}$ $=(\prod_{k=1}^{\infty})_{q}\left[\sum_{n_{k}=0}^{1}\exp_{q}\left(-\frac{(\varepsilon_{k}-\mu)n_{k}}{T}\right)\right]^{d}$ (A2)
$=(\prod_{k=1}^{\infty})_{q}\left[1+\exp_{q}\left(-\frac{(\varepsilon_{k}-\mu)}{T}\right)\right]^{d}.$
For non-interacting bosons, their grand-canonical partition function is written as
$Z_{q}^{B}=$ $(\prod_{k=1}^{\infty})_{q}\left[\sum_{n_{k}=0}^{\infty}\exp_{q}\left(-\frac{(\varepsilon_{k}-\mu)n_{k}}{T}\right)\right]^{d}$ (A3)
$=$ $(\prod_{k=1}^{\infty})_{q}\left[1+\cdots+{\rm exp}_{q}\left(-n_{k}\frac{\varepsilon_{k}-\mu}{T}\right)+\cdots\right]^{d}$
$\approx$ $(\prod_{k=1}^{\infty})_{q}\left[1-{\rm exp}_{q}\left(-\frac{\varepsilon_{k}-\mu}{T}\right)\right]^{-d},$
where the approximation $\textrm{exp}_{q}(x)\textrm{exp}_{q}(y)=\textrm{exp}_{q}(x+y+(1-q)xy)\approx\textrm{exp}_{q}(x+y)$ has been adopted [92].
Appendix B: The derivation of Eq. (14)
In the grand-canonical ensemble, the probability distribution of the state $R$ and the partition function in Eqs. (7) and (8) are, respectively, rewritten as
$p_{R}=\frac{1}{Z_{q}}\exp_{q}(-\alpha N_{R}-\beta E_{R}),$ (B1)
$Z_{R}=\sum_{R=1}^{W}\exp_{q}(-\alpha N_{R}-\beta E_{R}),$
where $\alpha\equiv-\mu/T$, $\beta\equiv 1/T$, $W$ is the total number of states.
The particle number $N$, the energy $E$, and the generalized force $Y$ are then, respectively, given by
$N$ $=\sum_{R}p_{R}^{q}N_{R}=\frac{\sum_{R}[\exp_{q}(-\alpha N_{R}-\beta E_{R})]^{q}N_{R}}{Z_{q}^{q}}$ (B2)
$=-\frac{\partial}{\partial\alpha}\ln_{q}Z_{q},$
$E$ $=\sum_{i}p_{R}^{q}E_{R}=\frac{\sum_{R}[\exp_{q}(-\alpha N_{R}-\beta E_{R})]^{q}E_{R}}{Z_{q}^{q}}$ (B3)
$=-\frac{\partial}{\partial\beta}\ln_{q}Z_{q},$
$Y$ $=\sum_{R}p_{R}^{q}\frac{\partial E_{R}}{\partial y}=\frac{\sum_{R}[\exp_{q}(-\alpha N_{R}-\beta E_{R})]^{q}\frac{\partial E_{R}}{\partial y}}{Z_{q}^{q}}$ (B4)
$=-\frac{1}{\beta}\frac{\partial}{\partial y}\ln_{q}Z_{q}.$
With Eqs. (B2), (B3) and (B4), the following relation is constructed,
$\beta\left(dE-Ydy+\frac{\alpha}{\beta}dN\right)$ (B5)
$=$ $-\beta d\left(\frac{\partial\ln_{q}Z_{q}}{\partial\beta}\right)+\frac{\partial\ln_{q}Z_{q}}{\partial y}dy-\alpha d\left(\frac{\partial\ln_{q}Z_{q}}{\partial\alpha}\right).$
Since $\ln_{q}Z_{q}$ is a function of $\alpha$, $\beta$, and $y$, its full differential is
$d(\ln_{q}Z_{q})=\frac{\partial\ln_{q}Z_{q}}{\partial\beta}d\beta+\frac{\partial\ln_{q}Z_{q}}{\partial\alpha}d\alpha+\frac{\partial\ln_{q}Z_{q}}{\partial y}dy.$ (B6)
Combing Eqs. (B5) and (B6), we can get
$\beta\left(dE-Ydy+\frac{\alpha}{\beta}dN\right)$ (B7)
$=$ $d\left(\ln_{q}Z_{q}-\alpha\frac{\partial\ln_{q}Z_{q}}{\partial\alpha}-\beta\frac{\partial\ln_{q}Z_{q}}{\partial\beta}\right).$
It has been shown that the following thermodynamic relation still holds in the non-extensive statistics [26, 20]
$dS=\frac{1}{T}(dE-Ydy-\mu dN).$ (B8)
Comparing Eq. (B7) with Eq. (B8), we obtain the entropy $S$ as
$S=\ln_{q}Z_{q}-\alpha\frac{\partial\ln_{q}Z_{q}}{\partial\alpha}-\beta\frac{\partial\ln_{q}Z_{q}}{\partial\beta}.$ (B9)
Therefore, in the non-extensive statistics the relation between the grand-canonical potential $\Omega$ and the grand-canonical partition function $Z_{q}$ is
$\Omega=$ $E-TS-\mu N$ (B10)
$=$ $-\frac{\partial}{\partial\beta}\ln_{q}Z_{q}-\mu\left(-\frac{\partial}{\partial\alpha}\ln_{q}Z_{q}\right)$
$-T\left(\ln_{q}Z_{q}-\alpha\frac{\partial\ln_{q}Z_{q}}{\partial\alpha}-\beta\frac{\partial\ln_{q}Z_{q}}{\partial\beta}\right)$
$=$ $-T\ln_{q}Z_{q}.$